A cylinder with a piston originally contains 5.0 L of He gas at a temperature of 310 K. When the gas is cooled the piston moves and the new volume is 2.4 L. What is the final temperature of the gas assuming that pressure and the number of moles of He remains constant?
1
Expert's answer
2020-03-27T08:58:56-0400

We know pV=nRT

As the number of moles and pressure remains constant we can write it as

V=kT

"\frac{V}{T}" =constant


"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{5}{310}=\frac{2.4}{T_2}"


"T_2"=148.8 K